# Extreme Precipitation Nowcasting using Multi-Task Latent Diffusion Models

### v2

@@ IV-D Evaluation based on Metrics
PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 MTLDM(oa)
1 0.0960 0.0866 0.1081 0.1141 0.0758 0.0923 0.0731 0.0968 0.1024 0.0638 0.0835 0.0624 0.0865 0.0924 0.0524
2 0.1101 0.1063 0.1209 0.1247 0.0921 0.1071 0.0924 0.1100 0.1137 0.0824 0.0988 0.0817 0.1007 0.1043 0.0719
3 0.1251 0.1160 0.1280 0.1308 0.1029 0.1230 0.1022 0.1174 0.1203 0.0944 0.1155 0.0916 0.1088 0.1114 0.0846
4 0.1354 0.1269 0.1342 0.1364 0.1111 0.1340 0.1133 0.1239 0.1262 0.1034 0.1270 0.1033 0.1158 0.1179 0.0941
5 0.1459 0.1360 0.1402 0.1419 0.1191 0.1445 0.1222 0.1296 0.1317 0.1120 0.1381 0.1128 0.1221 0.1237 0.1033
6 0.1557 0.1440 0.1465 0.1476 0.1262 0.1546 0.1300 0.1359 0.1375 0.1196 0.1487 0.1211 0.1287 0.1298 0.1113
7 0.1640 0.1506 0.1515 0.1521 0.1318 0.1630 0.1366 0.1408 0.1421 0.1257 0.1575 0.1282 0.1339 0.1347 0.1179
8 0.1721 0.1590 0.1571 0.1573 0.1391 0.1708 0.1444 0.1457 0.1471 0.1334 0.1657 0.1366 0.1392 0.1401 0.1259
9 0.1789 0.1645 0.1606 0.1608 0.1440 0.1777 0.1500 0.1494 0.1507 0.1386 0.1729 0.1425 0.1431 0.1439 0.1314
10 0.1848 0.1692 0.1641 0.1642 0.1485 0.1838 0.1549 0.1530 0.1543 0.1434 0.1793 0.1478 0.1469 0.1477 0.1365
11 0.1908 0.1744 0.1678 0.1679 0.1528 0.1896 0.1598 0.1566 0.1579 0.1479 0.1854 0.1531 0.1507 0.1516 0.1413
12 0.1961 0.1790 0.1712 0.1711 0.1571 0.1954 0.1645 0.1602 0.1613 0.1524 0.1914 0.1581 0.1544 0.1551 0.1461
13 0.2012 0.1836 0.1748 0.1745 0.1607 0.2001 0.1687 0.1634 0.1645 0.1562 0.1963 0.1627 0.1579 0.1586 0.1501
14 0.2039 0.1867 0.1770 0.1770 0.1637 0.2035 0.1722 0.1660 0.1672 0.1594 0.1999 0.1663 0.1606 0.1613 0.1535
15 0.2084 0.1907 0.1806 0.1803 0.1668 0.2078 0.1760 0.1693 0.1703 0.1627 0.2043 0.1703 0.1641 0.1646 0.1569
16 0.2123 0.1946 0.1850 0.1846 0.1700 0.2116 0.1795 0.1733 0.1744 0.1660 0.2084 0.1742 0.1683 0.1689 0.1604
@@
CRPS Analysis The CRPS describes the overall precipitation distribution, so we use the overall forecast radar image to evaluate it, that is, we calculate the index of the MTLDM (oa) (without any ambiguity, we will refer to it as the MTLDM). Fig.8 and Fig.9 show the CRPS values of the radar images at scales of 1, 4, and 16, and the detailed data of the comparative experiments are shown in Tables VI and VII. The image scaling is achieved through pooling operations. The experimental results consistently demonstrate the superiority of the MTLDM, which achieved the lowest CRPS values across all of the pooling scales and prediction ranges. This result highlights the effectiveness of the MTLDM in learning-rich temporal dependencies and generating accurate and well-calibrated probabilistic predictions. The other algorithms (especially the TRDMv2) managed to narrow their performance gaps with that of the MTLDM as the pooling scale increased. Moreover, we observed a general trend of increasing CRPS value over time for all of the algorithms, also demonstrating that prediction becomes increasingly difficult as time progresses. For a pooling scale of 1, the MTLDM’s CRPS values are significantly lower than those of the other algorithms at all time steps. Specifically, at the 80-minute prediction, the MTLDM’s CRPS value is about 0.17, approximately 7% lower than the second-best algorithm (TRDMv2), and about 20% lower than that of the worst-performing algorithm (PySTEPS). When the pooling scale increases to 4 and 16, although the CRPS values of all of the algorithms generally decrease, the gaps between them also narrow. Considering the time dimension, the CRPS values of all of the algorithms increase as the prediction step length increases, but the rate of increase gradually slows down. This indicates that the spatial shift of the degradation of the algorithm’s performance gradually decreases.
For a pooling scale of 1, the MTLDM’s CRPS values are significantly lower than those of the other four algorithms at all of the time steps. Specifically, for the 80-minute prediction (16 time steps), the MTLDM’s CRPS value is 0.1440, about 14% lower than that of the second-best algorithm DGMR(0.1679) and about 14% lower than that of the worst algorithm (PySTEPS; 0.1668). For the 5-minute prediction, the MTLDM’s advantage is even more pronounced, and its CRPS value (0.0842) is about 13% lower than that of the DGMR (0.0973) and about 21% lower than that of the PySTEPS (0.1062). For pooling scales of 4 and 16, the MTLDM’s CRPS values are significantly lower than those of the other algorithms at all of the time steps. The results demonstrate that the MTLDM achieves the lowest CRPS values at all of the time steps, and its advantage is particularly evident in the early stages of the prediction. For example, at the beginning of the prediction, the MTLDM’s CRPS values are 20% lower than those of the best of the other algorithms. Overall, the MTLDM demonstrates a clear advantage for this dataset, and its CRPS values are significantly lower than those of the other algorithms at all of the time steps, leading by about 10–20%. The DGMR and TRDM perform relatively similarly, while the TRDM and PySTEPS are slightly inferior. The performances of all of the algorithms decrease as the prediction time increases, but the rate of decrease slows down in the later stages.
@@ IV-E Evaluation based on Precipitation Events
Figure 12: Prediction performance and visualization evaluation for the overall precipitation events. (a) The precipitation event that occurred on October 15, 2021, at 13:25:00 in the SW dataset. (b)The precipitation event that occurred on February 28, 2022, at 11:42:00 in the MRMS dataset.
@@
In each dataset, we selected four precipitation events to evaluate the performance of the proposed prediction model. Two precipitation events were used to evaluate the prediction performance of the model for the overall precipitation (radar images are not segmented, see Figs  10-12 ), and the remaining two precipitation events were used to evaluate the prediction performance of segmented radar images at different precipitation intensities ($\geq$ 1 mm/h). Since precipitation of $\geq$ 4 mm/h is rare, we only make visual comparisons of the overall precipitation image and the precipitation forecast with a threshold of $\geq$ 1 mm/h (see Figs 13 and 14). During the evaluation process, the experimental results illustrate a comprehensive comparison of various precipitation forecasting models, benchmarked against the ground truth observations across different time steps, from T+5 minutes to T+65 minutes.
@@
Further analysis focusing on precipitation events with intensities of $\geq$ 1 mm/h reinforces these findings. MTLDM continues to exhibit remarkable accuracy in both spatial and temporal domains. Across both the subfigures (a) and (b) of Figs 13 and 14, MTLDM remains closely aligned with the ground truth, capturing the spread and intensity of precipitation with minimal deviation.MTLDM(oa) also has good prediction visual effects, but is slightly inferior to MTLDM in terms of the details of the predicted radar image. In contrast, other models, such as SSLDM, TRDMv1, TRDMv2, and DGMR, struggle to maintain this level of accuracy as the forecast window extends. While SSLDM and TRDMv1/v2 deliver relatively strong performance at shorter time intervals, they progressively lose accuracy at later time steps, with the predicted precipitation patterns becoming overly diffused or spatially shifted.
@@ V Conclusions
Accurate nowcasting of extreme precipitation is crucial for various weather-dependent decision-making processes, yet it remains a significant challenge. Our multi-task latent diffusion model (MTDM) offers a novel solution to this issue by addressing the limitations of current deep learning models in capturing spatial details in radar images, particularly in high-intensity precipitation areas. By decomposing radar images into sub-images representing different precipitation intensities and predicting each component separately, our approach achieves a superior performance across multiple evaluation metrics, based on a comparison with existing models, and thus can provide spatiotemporally consistent predictions for lead times at which existing methods often falter.
While the MTLDM has shown promising results, particularly in short-term prediction accuracy and consistency, challenges remain in predicting heavy precipitation at longer lead times. As revealed by our evaluations, including meteorological assessments, our approach offers substantial improvements over existing methods, yet the difficulty of accurately forecasting extreme precipitation persists across all approaches.
Furthermore, our work highlights the limitations of conventional verification metrics, which may not fully capture the operational utility of advanced models such as the MTLDM. This underscores the need to develop new quantitative measurements that better align with real-world applications, ensuring that improvements in predictive performance translate into practical benefits for decision-makers. We hope that our work will serve as a stepping stone for further integration in the fields of machine learning and environmental science, promoting the development of new data, methods, and evaluation techniques that enhance both the predictive accuracy and operational effectiveness in nowcasting of extreme weather events.

### v3

@@ IV-D Evaluation based on Metrics
PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa) PySTEPS DGMR TRDMv1 TRDMv2 DiffCast MTLDM(oa)
1 0.0960 0.0866 0.1081 0.1141 0.0725 0.0758 0.0923 0.0731 0.0968 0.1024 0.0605 0.0638 0.0835 0.0624 0.0865 0.0924 0.0508 0.0524
2 0.1101 0.1063 0.1209 0.1247 0.1033 0.0921 0.1071 0.0924 0.1100 0.1137 0.0951 0.0824 0.0988 0.0817 0.1007 0.1043 0.8644 0.0719
3 0.1251 0.1160 0.1280 0.1308 0.1225 0.1029 0.1230 0.1022 0.1174 0.1203 0.1155 0.0944 0.1155 0.0916 0.1088 0.1114 0.1073 0.0846
4 0.1354 0.1269 0.1342 0.1364 0.1365 0.1111 0.1340 0.1133 0.1239 0.1262 0.1303 0.1034 0.1270 0.1033 0.1158 0.1179 0.1227 0.0941
5 0.1459 0.1360 0.1402 0.1419 0.1526 0.1191 0.1445 0.1222 0.1296 0.1317 0.1470 0.1120 0.1381 0.1128 0.1221 0.1237 0.1399 0.1033
6 0.1557 0.1440 0.1465 0.1476 0.1605 0.1262 0.1546 0.1300 0.1359 0.1375 0.1552 0.1196 0.1487 0.1211 0.1287 0.1298 0.1487 0.1113
7 0.1640 0.1506 0.1515 0.1521 0.1658 0.1318 0.1630 0.1366 0.1408 0.1421 0.1611 0.1257 0.1575 0.1282 0.1339 0.1347 0.1552 0.1179
8 0.1721 0.1590 0.1571 0.1573 0.1756 0.1391 0.1708 0.1444 0.1457 0.1471 0.1711 0.1334 0.1657 0.1366 0.1392 0.1401 0.1654 0.1259
9 0.1789 0.1645 0.1606 0.1608 0.1820 0.1440 0.1777 0.1500 0.1494 0.1507 0.1779 0.1386 0.1729 0.1425 0.1431 0.1439 0.1727 0.1314
10 0.1848 0.1692 0.1641 0.1642 0.1903 0.1485 0.1838 0.1549 0.1530 0.1543 0.1864 0.1434 0.1793 0.1478 0.1469 0.1477 0.1814 0.1365
11 0.1908 0.1744 0.1678 0.1679 0.1943 0.1528 0.1896 0.1598 0.1566 0.1579 0.1905 0.1479 0.1854 0.1531 0.1507 0.1516 0.1858 0.1413
12 0.1961 0.1790 0.1712 0.1711 0.1971 0.1571 0.1954 0.1645 0.1602 0.1613 0.1953 0.1524 0.1914 0.1581 0.1544 0.1551 0.1892 0.1461
13 0.2012 0.1836 0.1748 0.1745 0.2032 0.1607 0.2001 0.1687 0.1634 0.1645 0.1997 0.1562 0.1963 0.1627 0.1579 0.1586 0.1954 0.1501
14 0.2039 0.1867 0.1770 0.1770 0.2075 0.1637 0.2035 0.1722 0.1660 0.1672 0.2042 0.1594 0.1999 0.1663 0.1606 0.1613 0.2003 0.1535
15 0.2084 0.1907 0.1806 0.1803 0.2139 0.1668 0.2078 0.1760 0.1693 0.1703 0.2106 0.1627 0.2043 0.1703 0.1641 0.1646 0.2068 0.1569
16 0.2123 0.1946 0.1850 0.1846 0.2159 0.1700 0.2116 0.1795 0.1733 0.1744 0.2128 0.1660 0.2084 0.1742 0.1683 0.1689 0.2091 0.1604
@@
CRPS Analysis The CRPS describes the overall precipitation distribution, so we use the overall forecast radar image to evaluate it, that is, we calculate the index of the MTLDM (oa) (without any ambiguity, we will refer to it as the MTLDM). Fig.7 and Fig.8 show the CRPS values of the radar images at scales of 1, 4, and 16, and the detailed data of the comparative experiments are shown in Tables VII and VIII. The image scaling is achieved through pooling operations. The experimental results consistently demonstrate the superiority of the MTLDM, which achieved the lowest CRPS values across all of the pooling scales and prediction ranges. This result highlights the effectiveness of the MTLDM in learning-rich temporal dependencies and generating accurate and well-calibrated probabilistic predictions. The other algorithms (especially the TRDMv2) managed to narrow their performance gaps with that of the MTLDM as the pooling scale increased. Moreover, we observed a general trend of increasing CRPS value over time for all of the algorithms, also demonstrating that prediction becomes increasingly difficult as time progresses. The CRPS indicators of the SW dataset at various scales are shown in Tables VII. For a pooling scale of 1, the MTLDM’s CRPS values are significantly lower than those of the other five algorithms at most of the time steps. Specifically, for the 80-minute prediction (16 time steps), the MTLDM’s CRPS value is 0.1440, about 10% lower than that of DiffCast (0.1600) and about 14% lower than that of DGMR (0.1679). For the 5-minute prediction, the MTLDM’s advantage is even more pronounced, with its CRPS value (0.0842) about 5% lower than that of DiffCast (0.0887), about 13% lower than that of DGMR (0.0973) and about 21% lower than that of PySTEPS (0.1062). For pooling scales of 4 and 16, the MTLDM’s CRPS values are significantly lower than those of the other algorithms at most time steps, particularly in the early stages. However, it’s worth noting that TRDMv1 outperforms MTLDM at later time steps (from step 9-15 at scale 1, and from steps 10-15 at scales 4 and 16). The results demonstrate that the MTLDM achieves the lowest CRPS values at the majority of time steps, with its advantage being particularly evident in the early stages of the prediction. At the beginning of the prediction, the MTLDM’s CRPS values are approximately 5-20% lower than those of the best competing algorithms. Overall, the MTLDM demonstrates a clear advantage for this dataset at early time steps, while TRDMv1 shows competitive performance at later time steps. DiffCast generally performs better than DGMR, TRDMv2, and PySTEPS across different scales, especially at early time steps. The performances of all algorithms decrease as the prediction time increases, but the rate of decrease slows down in the later stages.
The CRPS indicators of the MRMS dataset at various scales are shown in Tables VIII. For a pooling scale of 1 km, the performance comparison shows interesting patterns among the models. DiffCast achieves the best performance at the first time step with a CRPS of 0.0725, while MTLDM(oa) demonstrates superior performance across all subsequent time steps (2-16). As the pooling scale increases to 4 km and 16 km, all algorithms exhibit overall lower CRPS values, indicating improved prediction quality at coarser spatial resolutions. However, MTLDM(oa) consistently maintains the lowest CRPS values across all time steps at these larger scales, demonstrating its robust performance across different spatial resolutions. Among traditional methods, PySTEPS generally shows higher CRPS values than the deep learning approaches. DGMR performs better than TRDMv1 and TRDMv2 at early time steps, but this advantage diminishes in later prediction periods. DiffCast only outperforms MTLDM(oa) at the first time step at the 1 km scale, but falls behind at all other time steps and scales, showing a notable performance gap compared to MTLDM(oa).
@@
Figure 11: Prediction performance and visualization evaluation for the overall precipitation events. (a) The precipitation event that occurred on October 28, 2021, at 13:25:00 in the SW dataset. (b)The precipitation event that occurred on February 28, 2022, at 11:42:00 in the MRMS dataset.
@@ IV-E Evaluation based on Precipitation Events
In each dataset, we selected four precipitation events to evaluate the performance of the proposed prediction model. Two precipitation events were used to evaluate the prediction performance of the model for the overall precipitation (radar images are not segmented, see Figs  9-11 ), and the remaining two precipitation events were used to evaluate the prediction performance of segmented radar images at different precipitation intensities ($\geq$ 1 mm/h). Since precipitation of $\geq$ 4 mm/h is rare, we only make visual comparisons of the overall precipitation image and the precipitation forecast with a threshold of $\geq$ 1 mm/h (see Figs 12 and 13). During the evaluation process, the experimental results illustrate a comprehensive comparison of various precipitation forecasting models, benchmarked against the ground truth observations across different time steps, from T+5 minutes to T+65 minutes.
@@
Further analysis focusing on precipitation events with intensities of $\geq$ 1 mm/h reinforces these findings. MTLDM continues to exhibit remarkable accuracy in both spatial and temporal domains. Across both the subfigures (a) and (b) of Figs 11 and 12, MTLDM remains closely aligned with the ground truth, capturing the spread and intensity of precipitation with minimal deviation.MTLDM(oa) also has good prediction visual effects, but is slightly inferior to MTLDM in terms of the details of the predicted radar image. In contrast, other models, such as SSLDM, TRDMv1, TRDMv2, and DGMR, struggle to maintain this level of accuracy as the forecast window extends. While SSLDM and TRDMv1/v2 deliver relatively strong performance at shorter time intervals, they progressively lose accuracy at later time steps, with the predicted precipitation patterns becoming overly diffused or spatially shifted. Due to limited network capacity, the precipitation results predicted by Diffcast show a trend of gradually weakening over time, resulting in significant differences between the long-term predicted precipitation areas and the actual observed values. This model limitation is clearly reflected in Figures 11.a and 12.a, where the deviation between the predicted values and the actual precipitation distribution increases significantly with the extension of the prediction time.
@@ V Conclusions
Accurately predicting extreme precipitation is critical for a variety of weather-dependent decision-making processes, but it remains a major challenge. Our multi-task latent diffusion model (MTLDM) provides an innovative solution to this problem by addressing the limitations of current deep learning models in capturing spatial details in radar images, especially in areas of high-intensity precipitation. MTLDM decomposes radar images into sub-images representing different precipitation intensities and predicts each component separately, thereby achieving better performance than existing models on multiple evaluation metrics and providing spatiotemporally consistent forecasts over forecast periods where existing methods often fail. An important feature of MTLDM is the ability to decode the latent space. The main advantage of this approach is that it can directly operate on the latent space using a trained rainfall model and generate forecasts without additional training. However, this strategy of relying on the latent space also brings limitations: model performance is largely limited by the quality and accuracy of the latent space representation. If the latent space encoding is not sufficient to capture complex meteorological patterns, especially nonlinear changes in extreme weather events, it will limit the overall forecasting ability. While MTLDM performs well in terms of accuracy and consistency in short-term forecasts, it remains challenging to predict heavy precipitation over longer forecast periods. Our evaluation shows that although MTLDM is a significant improvement over existing methods, the difficulty of accurately predicting extreme precipitation remains in all methods, and the model effect gradually decreases as the forecast time horizon increases. In addition, our study also highlights the limitations of conventional validation indicators, which may not fully reflect the practical application value of advanced models such as MTLDM. This suggests that new quantitative evaluation methods need to be developed to better adapt them to practical applications and ensure that the improvement in prediction performance can be translated into practical benefits for decision makers. We hope that this study will promote further integration of machine learning and environmental science, promote the development of new data, methods and evaluation techniques, and improve the prediction accuracy and practical utility of extreme weather event nowcasting.